Problem A.1: Interstellar Mission (4 Points)

approximated density of

(a) Calculate the (b)

1

You are on an interstellar mission from the Earth to the 8.7 light-years distant star Sirius. Your spaceship can travel with 70% the speed of light and has a cylindrical shape with a diameter of 6 m at the front surface and a length of 25 m. You have to cross the interstellar medium with an hydrogen atom/m

time it takes

your spaceship to reach Sirius.

Determine the mass of interstellar gas that collides with your spaceship during the mission.

Note: Use 1.673 x 10-27 kg as proton mass.
1
Expert's answer
2020-06-23T13:00:27-0400

Solution.

Earth to Sirius"=8.7"  light-years;

"70\%" of c;

"d=6m;"

"l=25m;"

"m_p=1.673\sdot10^{-27} kg ;"

a)"t=\dfrac{d}{v};"


"t=\dfrac{8.7}{0.7}=12.43 years;"


b) "S=\pi r^2+2\pi rl;"

"S=\pi\sdot (3m)^2+2\sdot\pi\sdot3m\sdot25m=159\pi m^2=499.26m^2;"

"h=8.7\sdot9.46\sdot10^{15}m=8.23\sdot10^{16}m;"

"V=499.26m^2\sdot8.23\sdot10^{16}m=4.108\sdot10^{19}m^3;"

m="D\sdot V;"

"m=1.673\sdot10^{-27}kg/m^3\sdot4.108\sdot10^{19}m^3=6.87\sdot10^{-8}kg;"

Answer: "a) t=12.43years;"

"b)m=6.87\sdot10^{-8}kg."